# Holmes: Towards Effective and Harmless Model Ownership Verification to Personalized Large Vision Models via Decoupling Common Features

### v1

@@ IV-B Training Ownership Meta-Classifier
As we described in Section IV-A, we utilize the output difference between $V$ and $P$ on $\mathcal{D}_{s}$ as the representation of $V$’s dataset-specific features. Since the output difference between $V$ and $P$ is a distribution rather than a single value, we need to train an additional binary meta-classifier $C:\mathbb{R}^{|\bm{w}|}\rightarrow\{-1,+1\}$ to determine whether the suspicious model contains the knowledge of the dataset-specific features.
@@
The key insight is to contrast two types of output differences: (1) the output difference between $V$ and $P$, representing unique dataset-specific features of $V$; (2) the output difference between $B$ and $P$, denoting features distinct from $V$’ dataset-specific features, as $B$ shares common features with $V$ but learns different dataset-specific features from $\mathcal{D}_{f}$. By comparing these differences, $C$ learns to identify dataset-specific features of $V$ from irrelevant features. Formally, we define the output differences of a data sample $\bm{x}$ as:
@@
where $\mathcal{L}(\cdot)$ is the loss function ($i.e.$, cross-entropy), $\bm{w}$ represents the parameters of the meta-classifier $C$, and $l\in\{+1,-1\}$ denotes the label indicating whether the input corresponds to the dataset-specific features ($i.e.$, ‘$l=+1$’) learned by victim model or unrelated features ($i.e.$, ‘$l=-1$’).
IV-C Ownership Verification with Hypothesis Test
In the verification process, given a suspicious model $S$, the defender can examine it via the meta-classifier output $C(\bm{O}_{s}(\bm{x}))$ for samples $(\bm{x},y)\in\mathcal{D}_{s}$, where $\bm{O}_{s}(\bm{x})=S(\bm{x})-P(\bm{x})$. If $C(\bm{O}_{s}(\bm{x}))=+1$, the suspicious model is considered stolen from the victim. However, this decision is susceptible to the randomness of the selected $\bm{x}$. To mitigate this issue, we adopt a hypothesis test-based method inspired by existing works [25, 50, 24], as described below:
@@ IV-C Ownership Verification with Hypothesis Test
Let $\bm{X}$ be a random variable representing samples from $\mathcal{D}_{s}$. Define $\mu_{S}$ and $\mu_{B}$ as the posterior probabilities of the events $C(\bm{O}_{s}(\bm{x}))=+1$ and $C(\bm{O}_{s}(\bm{x}))=-1$, respectively. Given the null hypothesis $H_{0}:\mu_{B}+\tau=\mu_{S}$ ($H_{1}:\mu_{B}+\tau<\mu_{S}$) where $\tau\in[0,1]$, we claim that the suspicious model $S$ is stolen from the victim (with $\tau$certainty) if and only if $H_{0}$ is rejected.
In practice, we randomly sample $m$ samples from $\mathcal{D}_{s}$ to conduct the pair-wise T-test [58] and calculate its p-value. When the p-value is smaller than the significance level $\alpha$, $H_{0}$ is rejected. Besides, we also calculate the confidence score $\Delta\mu=\mu_{S}-\mu_{B}$ to represent the verification confidence. The larger the $\Delta\mu$, the more confident the verification.
IV-D Robustness Analysis of Holmes’s Ownership Verification
Since perfect accuracy (100%) is usually unachievable for $C$, it raises concerns about the reliability of our ownership verification under classifier imperfections. In this section, we demonstrate that our verification framework remains statistically robust when $C$ exhibits realistic error rates. The success condition of verification is formalized as follows.
@@ IV-D Robustness Analysis of Holmes’s Ownership Verification
where $\Delta=t_{1-\alpha}^{4}+4t_{1-\alpha}^{2}(m-1)(\beta+\tau)(1-\beta-\tau)$, $t_{1-\alpha}$ is the $(1-\alpha)$-quantile of $t$-distribution with $(m-1)$ degrees of freedom, and m is the sample size of $\bm{X}$.
In general, Theorem 1 indicates two critical properties: (1) Our hypothesis test-based ownership verification can succeed if the identification success rate of $C$ is higher than a threshold (which is not necessarily 100%). (2) The defender can claim the model ownership with limited queries to the suspicious model $S$ if the identification success rate is high enough. The proof of Theorem 1 is included in our appendix.
V Experiments
((a))
((b))
((c))
((d))
((e))
((f))
((g))
((h))
((i))
((j))
@@ V-A Main Settings
Dataset Selection. We evaluate the effectiveness of our method using the CIFAR-10 [54] and ImageNet [56] datasets. CIFAR-10 consists of 60,000 images across 10 categories, with 50,000 samples allocated for training and 10,000 for testing. For ImageNet, we use a subset of 200 randomly selected classes to balance computational efficiency and representativeness. Each class in this subset includes 5,000 samples for training set and 50 samples for testing set.
@@
Settings of Personalization. We adopt the CLIP [2] model (having a ViT-B/32 Transformer architecture as an image encoder) as the victim model for experiments on the CIFAR-10 and ImageNet datasets. Specifically, we use the pre-trained CLIP model from a public repository [^1] https://huggingface.co/openai/clip-vit-base-patch32 as the the foundation model. For personalization, the victim models are obtained by fine-tuning the pre-trained foundation model on CIFAR-10 and ImageNet for 20 epochs. The learning rate is set to $1\times 10^{-5}$. To improve fine-tuning effectiveness, the learning rate for the linear layers ($i.e.$, the final layers of CLIP model) is increased by a factor of 10, reaching $1\times 10^{-4}$.
@@
Settings of Model Stealing. Following the settings in [25, 50, 24], we perform the model stealing attacks described in Section II-B to assess the effectiveness of the proposed method. In particular, (1) Direct-copy: The adversary directly copies the victim model. (2) Fine-tuning: The adversary obtains stolen models by fine-tuning victim models for 5 epochs on different substitute datasets. Following the settings in [25, 50, 24], we randomly select 500,000 images from the original TinyImages dataset [59] to construct the substitute dataset for the experiments on CIFAR-10 dataset. For ImageNet experiments, we randomly sample 200 classes distinct from those used to fine-tune the foundation model to generate the substitute dataset. (3) Distillation: In particular, we implement the knowledge distillation method [11] using open-sourced code [^2] https://github.com/zhengli97/ctkd. The stolen model is trained for 100 epochs using the SGD optimizer with an initial learning rate of 0.1, momentum of 0.9, and a weight decay of $1\times 10^{-4}$. (4) Data-free knowledge distillation (DFKD): We adopt NAYER [44] as the stealing attack method for DFKD. The NAYER implementation is based on the open-source code [^3] https://github.com/tmtuan1307/NAYER. For both CIFAR-10 and ImageNet datasets, the stolen models are trained for 300 epochs. Other experimental settings follow those described in the original NAYER work. (5) Label-query: The stolen model is trained for 200 epochs on substitute datasets (identical to those used in the fine-tuning attack) using labels generated by the victim model. (6) Logit-query: For this attack, the stolen model is trained using logit-based knowledge distillation [13] with its open-sourced code [^4] https://github.com/sunshangquan/logit-standardization-KD with the same substitute datasets as in the fine-tuning attack. This stolen model is trained for 100 epochs with SGD optimizer and an initial learning rate of 0.1, momentum of 0.9, and weight decay of $1\times 10^{-4}$. Besides, we also conduct experiments of examining a suspicious model which is not stolen from the victim (dubbed ‘Independent’) for reference. To construct the independent baseline models, we utilize a pre-trained CLIP model with a ViT-B/16 Transformer architecture as the image encoder, sourced from a public repository [^5] https://huggingface.co/openai/clip-vit-base-patch16. This foundation model is fine-tuned on TinyImages and the 200-class substitute dataset from ImageNet to create independent models for CIFAR-10 and ImageNet experiments, respectively. The fine-tuning settings for these independent models are consistent with those used for the victim models.
@@
Settings of Defenses. We compare our method with model watermarking methods and model fingerprinting methods. For model watermarking baselines, we examine methods including BadNets watermarking [28] implemented via BackdoorBox [60], gradient matching (GM) [16], entangled watermarks (EWE) [17], PTYNet [18], UAE [19], and MOVE [24]. All watermarking methods employ a consistent 10% modification rate on training samples, with implementation details following their official open-source repositories. Figure 3 illustrates representative examples of modified images across different defenses, including BadNets’ black square trigger, poisoned image in GM, EWE’s out-of-distribution samples from SVHN dataset [61], PTYNet’s semantic background triggers, and the oil painting style image of MOVE. Specifically, MOVE adopts different approaches for white-box and black-box settings. Since this paper focuses on verification in black-box settings, all experiments are conducted under black-box settings.
@@
For model fingerprinting baselines, we compare our method against dataset inference (DI) [25], universal adversarial perturbations (UAPs) [26], and Metafinger [27]. DI adopts different approaches for white-box and black-box settings. Following the black-box verification focus of this work, DI is evaluated under its black-box configurations.
@@
For our method, we construct the poisoned model $P$ by fine-tuning $V$ on $\mathcal{D}_{p}\cup\mathcal{D}_{b}$ ($\mathcal{D}_{b}\triangleq\mathcal{D}\backslash\mathcal{D}_{s}$) for 5 epochs. For both CIFAR-10 and ImageNet, we set the modification rate $\gamma\%=10\%$, resulting in $|\mathcal{D}_{s}|=5,000$ (CIFAR-10) and $|\mathcal{D}_{s}|=1,000$ (ImageNet). The benign shadow model $B$ is obtained by fine-tuning the foundation model $F$ on $\mathcal{D}_{f}$ for 20 epochs with a dropping rate $\lambda\%=50\%$ for both datasets. Both $P$ and $B$ use a base learning rate of $1\times 10^{-5}$. For $B$, the learning rate for the final linear layers increases to $1\times 10^{-4}$ to enhance fine-tuning effectiveness. The meta-classifier $C$ is trained for 300 epochs with a learning rate of 0.01. The hypothesis test is conducted at significance level $\alpha=0.01$.
@@
TABLE IV: Results of defenses on the CIFAR-10 dataset. We hereby mark failed verification cases ($i.e.$, p-value $>0.01$ for stolen models and p-value $<0.01$ for independent models) in red.
@@
Holmes 0.18 $10^{-14}$ 0.12 $10^{-9}$ 0.18 $10^{-12}$ 0.13 $\mathbf{10^{-9}}$ 0.05 $\mathbf{10^{-4}}$ 0.14 $\mathbf{10^{-8}}$0.03 0.20
TABLE V: Results of defenses on the ImageNet dataset. We hereby mark failed verification cases ($i.e.$, p-value $>0.01$ for stolen models and p-value $<0.01$ for independent models) in red.
Model Stealing $\rightarrow$ Direct-copy Fine-tuning Distillation DFKD Label-query Logit-query Independent
@@
EWE $-10^{-5}$ 0.99 -0.02 0.99 $-10^{-3}$ 0.94 -0.07 0.99 $-10^{-3}$ 0.98 -0.01 0.98 -0.06 0.99
PTYNet 0.11 $10^{-3}$0.02 0.08 -0.02 0.84 -0.03 0.96 -0.03 0.75 -0.01 0.03 -0.01 0.52
UAE 0.35 $10^{-3}$ $10^{-4}$ 0.26 $10^{-4}$ 0.22 $10^{-5}$ 0.29 $10^{-7}$ 0.21 $-10^{-7}$ 0.33 $10^{-5}$ 0.23
MOVE $10^{-4}$ $10^{-4}$ $10^{-4}$ $10^{-5}$ $10^{-4}$ 0.05 0.11 $10^{-3}$ $-10^{-4}$ 0.44 $-10^{-4}$ 0.56 0.00 0.99
Fingerprint UAPs 0.52 $10^{-7}$ 0.01 $10^{-3}$ $10^{-4}$ 0.27 -0.01 0.13 -0.10 0.33 -0.12 0.24 -0.01 0.21
@@
Holmes 0.37 $\mathbf{10^{-36}}$ 0.17 $\mathbf{10^{-10}}$ 0.27 $\mathbf{10^{-20}}$ 0.10 $\mathbf{10^{-4}}$ 0.37 $\mathbf{10^{-32}}$ 0.22 $\mathbf{10^{-11}}$0.61 0.99
Evaluation Metric. We use the confidence score $\Delta\mu$ and p-value as the metric for our evaluation. Both of them are calculated based on the hypothesis test with 100 sampled images, $i.e.$, $m=100$. For cases involving stolen models, a lower p-value and a higher $\Delta\mu$ indicate stronger verification performance. Conversely, for independent models, a higher p-value and a lower $\Delta\mu$ suggest better resistance to misclassification. The best results among all defenses are highlighted in bold, while failed verifications ($i.e.$, p-value $>\alpha$ for stolen models and p-value $<\alpha$ for independent models, where $\alpha=0.01$) are marked in red.
We note that some baseline methods ($e.g.$, model watermarking with BadNets) do not use hypothesis test for ownership verification. These methods typically determine ownership by assessing whether the probability assigned by the suspicious model to a watermarked sample for a specific target label is significantly higher than that assigned to its original version. For fair comparison, we also adopt hypothesis test for these methods as follows.
@@
Let $\bm{X}^{\prime}$ and $\bm{X}$ denote the variable of the watermarked sample and that its benign version, respectively. Define $\mu_{S}$ and $\mu_{B}$ as $\mu_{S}\triangleq S(\bm{X}^{\prime})_{y_{t}}$ and $\mu_{B}\triangleq S(\bm{X})_{y_{t}}$, where $y_{t}$ is the target label. Given the null hypothesis $H_{0}:\mu_{B}+\tau=\mu_{S}$ ($H_{1}:\mu_{B}+\tau<\mu_{S}$) where $\tau\in[0,1]$, we claim that the suspicious model $S$ is stolen from the victim (with $\tau$certainty) if and only if $H_{0}$ is rejected.
Specifically, the $\Delta\mu$ is calculated as $\Delta\mu=\mu_{S}-\mu_{B}$.
V-B Main Results
As shown in Tables IV-V, our method achieves the best model verification performance in almost all cases. Notably, it is the only method that achieves reliable verification across all model stealing scenarios while avoiding any misjudgments for the independent models. For example, in the logit-query attack on the CIFAR-10 dataset, our method achieves a p-value of $10^{-8}$, significantly outperforming MOVE, the only baseline capable of successful verification in this scenario, which achieves a p-value of 0.01. In the cases of direct-copy, distillation, and fine-tuning, methods such as BadNets achieve competitive results. Nevertheless, our method can still easily make correct predictions in these cases. Besides, on the ImageNet dataset, our method demonstrates superior performance across all scenarios, including both model stealing cases and the independent model case. For example, in the distillation scenario, our method achieves a p-value of $10^{-20}$, outperforming DI’s $10^{-8}$, while all other baseline methods fail to provide successful model ownership verification. Additionally, in the independent model scenario for both CIFAR-10 and ImageNet datasets, our method consistently avoids misjudgments, while DI misclassifies the benign independent models as stolen models. This indicates that our method effectively decouples common features while emphasizing dataset-specific features, ensuring reliable model ownership verification across diverse datasets and model stealing attack methods.
V-C Effects of Key Hyper-parameters

### v2

@@ 4.2 Training Meta-Classifier
As we described in Section 4.1, we utilize the output difference between $V$ and $P$ on $\mathcal{D}_{s}$ as the representation of $V$’s dataset-specific features. Since the output difference between $V$ and $P$ is a distribution rather than a single value, we need to train an additional binary meta-classifier $C:\mathbb{R}^{|\bm{w}|}\rightarrow\{-1,+1\}$ ($\bm{w}$ represents the parameters of the meta-classifier $C$) to determine whether the suspicious model contains the knowledge of the dataset-specific features.
@@
The key insight is to contrast two types of output differences: (1) the output difference between $V$ and $P$, representing unique dataset-specific features of $V$; (2) the output difference between $B$ and $P$, denoting features distinct from $V$’s dataset-specific features, as $B$ shares common features with $V$ but learns different dataset-specific features from $\mathcal{D}_{f}$. By comparing these differences, $C$ learns to identify dataset-specific features of $V$ from irrelevant features. Formally, we define the output differences of a data sample $\bm{x}$ as:
@@
where $\mathcal{L}(\cdot)$ is the loss function ($i.e.$, cross-entropy), $\bm{w}$ represents the parameters of the meta-classifier $C$, and $l\in\{+1,-1\}$ denotes the label indicating whether the input corresponds to the dataset-specific features ($i.e.$, ‘$l=+1$’) learned by victim model or unrelated features ($i.e.$, ‘l=1l=-1’).
4.3 Ownership Verification with Hypothesis Test
In the verification process, given a suspicious model $S$, the defender can examine it via the meta-classifier output $C(\bm{O}_{s}(\bm{x}))$ for samples $(\bm{x},y)\in\mathcal{D}_{s}$, where $\bm{O}_{s}(\bm{x})=S(\bm{x})-P(\bm{x})$. If $C(\bm{O}_{s}(\bm{x}))=+1$, $S$ is considered stolen from the victim. However, this decision is susceptible to the randomness of the selected $\bm{x}$. To mitigate this issue, we adopt a hypothesis test-based method inspired by existing works [maini2021dataset, li2022defending, li2025move], as described below:
@@ 4.3 Ownership Verification with Hypothesis Test
Let $\bm{X}$ be a random variable representing samples from $\mathcal{D}_{s}$. Define $\mu_{S}$ and $\mu_{B}$ as the posterior probabilities of the events $C(\bm{O}_{s}(\bm{x}))=+1$ and $C(\bm{O}_{s}(\bm{x}))=-1$, respectively. Given the null hypothesis $H_{0}:\mu_{B}+\tau=\mu_{S}$ ($H_{1}:\mu_{B}+\tau<\mu_{S}$) where $\tau\in[0,1]$, we claim that the suspicious model $S$ is stolen from the victim (with $\tau$-certainty) if and only if $H_{0}$ is rejected.
In practice, we randomly sample $m$ samples from $\mathcal{D}_{s}$ to conduct the pair-wise T-test [hogg2005introduction] and calculate its corresponding p-value. When the p-value is smaller than the pre-defined significance level $\alpha$, $H_{0}$ is rejected. Besides, we also calculate the confidence score $\Delta\mu=\mu_{S}-\mu_{B}$ to represent the verification confidence. The larger the $\Delta\mu$, the more confident the verification.
4.4 Robustness Analysis of Holmes’s Ownership Verification
Since perfect accuracy (100%) is usually unachievable for the meta-classifier $C$, it raises concerns about the reliability of our ownership verification under classifier imperfections. In this section, we demonstrate that our verification framework remains statistically robust when $C$ exhibits realistic error rates. The success condition of verification is formalized as follows.
@@ 4.4 Robustness Analysis of Holmes’s Ownership Verification
where $\Delta=t_{1-\alpha}^{4}+4t_{1-\alpha}^{2}(m-1)(\beta+\tau)(1-\beta-\tau)$, $t_{1-\alpha}$ is the $(1-\alpha)$-quantile of $t$-distribution with (m1)(m-1) degrees of freedom, and m is the sample size of $\bm{X}$.
In general, Theorem 1 indicates two critical properties: (1) Our hypothesis test-based ownership verification can succeed if the identification success rate of $C$ is higher than a threshold (which is not necessarily 100%). (2) The defender can claim the model ownership with limited queries to the suspicious model $S$ if the identification success rate is high enough. The detailed proof of Theorem 1 is included in Appendix A.
5 Experiments
@@ 5.1 Main Settings
Dataset Selection. We evaluate the effectiveness of Holmes using the CIFAR-10 [krizhevsky2009learning] and ImageNet [deng2009imagenet] datasets. CIFAR-10 consists of 60,000 images across 10 categories, with 50,000 samples allocated for training and 10,000 for testing. For ImageNet, we use a subset of 200 randomly selected classes to balance computational efficiency and representativeness. Each class in this subset includes 2,500 randomly selected samples for the training set and 50 samples for the testing set.
@@
Settings of Personalization. We adopt the CLIP [radford2021learning] model (with a ViT-B/32 Transformer architecture [vaswani2017attention] as the image encoder) as the victim model for experiments on the CIFAR-10 and ImageNet datasets. Specifically, we use the pre-trained CLIP model from the public repository [^1] https://huggingface.co/openai/clip-vit-base-patch32 as the foundation model. For personalization, the victim models are obtained by fine-tuning the pre-trained foundation model on CIFAR-10 and ImageNet datasets for 20 epochs, respectively. The learning rate is set to $1\times 10^{-5}$. To improve fine-tuning effectiveness, the learning rate for the linear layers ($i.e.$, the final layers of CLIP models) is increased by a factor of 10, reaching $1\times 10^{-4}$.
Settings of Model Stealing. Following the settings in [maini2021dataset, li2022defending, li2025move], we perform the model stealing attacks described in Section 2.2 to assess the effectiveness of the proposed method.
Before detailing each attack, we first clarify the architectures and datasets used for the stolen models across different scenarios. For Direct-copy and Fine-tuning attacks, the stolen model adopts the same architecture as the victim model. For all distillation-based attacks, including distillation with access to the training dataset (Data Distill), data-free distillation (Data-free Distill), Hard Distill, and Soft Distill, we employ a ResNet-18 network as the stolen model.
Regarding the datasets used in model stealing, we construct substitute datasets for Fine-tuning, Hard Distill, and Soft Distill attacks, following [maini2021dataset, li2022defending, li2025move]. For CIFAR-10 experiments, we randomly sample 500,000 images from the TinyImages dataset [birhane2021large] as the substitute dataset. For ImageNet experiments, we select 2,500 samples from the same 200 classes used to fine-tune the victim model, ensuring that all selected samples are disjoint from those used during personalization.
@@
In particular, (1) Direct-copy: The adversary directly copies the victim model. (2) Fine-tuning: The adversary obtains stolen models by fine-tuning victim models for 5 epochs on substitute datasets. (3) Data Distill: We implement the CTKD method [li2023curriculum] using open-source code [^2] https://github.com/zhengli97/ctkd. The stolen model is trained for 100 epochs on the training set of $V$ using the SGD optimizer with an initial learning rate of 0.1, momentum of 0.9, and a weight decay of $1\times 10^{-4}$. (4) Data-free Distill: We adopt NAYER [tran2024nayer] as the attack method, using its open-source code [^3] https://github.com/tmtuan1307/NAYER. For both datasets, the stolen model is trained for 300 epochs, with other settings following the original work. (5) Hard Distill: The stolen model is trained for 200 epochs on the substitute datasets using hard labels generated by the victim model. We adopt the SGD optimizer with an initial learning rate of 0.1, momentum of 0.9, and a weight decay of $1\times 10^{-4}$. (6) Soft Distill: The stolen model is trained using logit-based knowledge distillation [sun2024logit], implemented with the open-source code [^4] https://github.com/sunshangquan/logit-standardization-KD. Training is performed for 100 epochs using the SGD optimizer with an initial learning rate of 0.1, momentum of 0.9, and a weight decay of $1\times 10^{-4}$. Besides, we conduct experiments of examining a suspicious model which is not stolen from the victim (dubbed ‘Independent’) for reference. We utilize a pre-trained CLIP model with a ViT-B/16 Transformer architecture as the image encoder, sourced from the public repository [^5] https://huggingface.co/openai/clip-vit-base-patch16. This foundation model is fine-tuned on substitute datasets to create independent models for CIFAR-10 and ImageNet experiments, respectively. The fine-tuning settings for these independent models are consistent with those used for the victim models.
@@
Settings of Defenses. We compare Holmes with model watermarking and model fingerprinting methods. For model watermarking baselines, we examine methods including BadNets watermarking [gu2019badnets] implemented via BackdoorBox [li2022backdoorbox], gradient matching (GM) [geiping2021witches], entangled watermarks (EWE) [jia2021entangled], PTYNet [wang2023free], UAE [zhu2024reliable], and MOVE [li2025move]. All watermarking methods employ a consistent 10% modification rate on training samples, with implementation details following their official open-source repositories. Figure 3 illustrates representative examples of modified images across different defenses, including BadNets’ black square trigger, poisoned image in GM, EWE’s out-of-distribution samples from SVHN dataset [netzer2011reading], PTYNet’s semantic background triggers, and the oil painting style image of MOVE. Specifically, MOVE adopts different approaches for white-box and black-box settings. Since this paper focuses on verification in black-box settings, all experiments are conducted under black-box settings.
@@
For model fingerprinting baselines, we compare Holmes against dataset inference (DI) [maini2021dataset], universal adversarial perturbations (UAPs) [peng2022fingerprinting], and Metafinger [yang2022metafinger]. DI adopts different approaches for white-box and black-box settings. Following the black-box verification focus of this work, DI is evaluated under its black-box configurations.
@@
For our method, we construct the poisoned model $P$ by fine-tuning $V$ on $\mathcal{D}_{p}\cup\mathcal{D}_{b}$ ($\mathcal{D}_{b}\triangleq\mathcal{D}\backslash\mathcal{D}_{s}$) for 5 epochs. For both CIFAR-10 and ImageNet, we set the modification rate $\gamma\%=10\%$, resulting in $|\mathcal{D}_{s}|=5,000$ (CIFAR-10) and $|\mathcal{D}_{s}|=500$ (ImageNet). The poisoned trigger is a black square in the top-left corner, as shown in Figure 3j. The target label is set to “automobile” for CIFAR-10 and “white shark” for ImageNet. The benign shadow model $B$ is obtained by fine-tuning the foundation model $F$ on $\mathcal{D}_{f}$ for 20 epochs with a dropping rate $\lambda\%=50\%$ for both datasets. Both $P$ and $B$ use a base learning rate of $1\times 10^{-5}$. For $B$, the learning rate for the final linear layers increases to $1\times 10^{-4}$ to enhance fine-tuning effectiveness. The meta-classifier $C$ is trained for 300 epochs with a learning rate of 0.01. The hypothesis test is conducted at significance level $\alpha=0.01$.
@@
Table 4: Results of defenses on the CIFAR-10 dataset. We mark failed verification cases ($i.e.$, p-value >0.01>0.01 for stolen models and p-value <0.01<0.01 for independent models) in red.
@@
Holmes 0.18 $10^{-14}$ 0.12 $10^{-9}$ 0.18 $10^{-12}$ 0.13 $\mathbf{10^{-9}}$ 0.05 $\mathbf{10^{-4}}$ 0.14 $\mathbf{10^{-8}}$ -0.03 0.20
Table 5: Results of defenses on the ImageNet dataset. We mark failed verification cases ($i.e.$, p-value >0.01>0.01 for stolen models and p-value <0.01<0.01 for independent models) in red.
Model Stealing $\rightarrow$ Direct-copy Fine-tuning Data Distill Data-free Distill Hard Distill Soft Distill Independent
@@
EWE $10^{-5}$ 0.99 -0.02 0.99 $10^{-3}$ 0.94 -0.07 0.99 $10^{-3}$ 0.98 -0.01 0.98 -0.06 0.99
PTYNet 0.11 $10^{-3}$ -0.02 0.08 -0.02 0.84 -0.03 0.96 -0.03 0.75 -0.01 0.03 -0.01 0.52
UAE 0.35 $10^{-3}$ $10^{-4}$ 0.26 $10^{-4}$ 0.22 $10^{-5}$ 0.29 $10^{-7}$ 0.21 $10^{-7}$ 0.33 $10^{-5}$ 0.23
MOVE $10^{-4}$ $10^{-4}$ $10^{-4}$ $10^{-5}$ $10^{-4}$ 0.05 0.11 $10^{-3}$ $10^{-4}$ 0.44 $10^{-4}$ 0.56 0.00 0.99
[…]
@@
Holmes 0.37 $\mathbf{10^{-36}}$ 0.17 $\mathbf{10^{-10}}$ 0.27 $\mathbf{10^{-20}}$ 0.10 $\mathbf{10^{-4}}$ 0.37 $\mathbf{10^{-32}}$ 0.22 $\mathbf{10^{-11}}$ -0.61 0.99
Evaluation Metric. We use the confidence score $\Delta\mu$ and p-value as the metric for our evaluation. Both of them are calculated based on the hypothesis test with 100 sampled images, $i.e.$, $m=100$. For cases involving stolen models, a lower p-value and a higher $\Delta\mu$ indicate stronger verification performance. Conversely, for independent models, a higher p-value and a lower $\Delta\mu$ suggest better resistance to misclassification. The best results among all defenses are highlighted in bold, while failed verifications ($i.e.$, p-value $\alpha$ for stolen models and p-value $\alpha$ for independent models, where $\alpha=0.01$) are marked in red.
We note that some baseline methods ($e.g.$, model watermarking with BadNets) do not use hypothesis test for model ownership verification. These methods typically determine model ownership by assessing whether the probability assigned by the suspicious model to a watermarked sample for a specific target label is significantly higher than that assigned to its original version. For fair comparison, we also adopt hypothesis test for these methods, as follows:
@@
Let $\bm{X}^{\prime}$ and $\bm{X}$ denote the variable of the watermarked sample and its benign version, respectively. Define $\mu_{S}$ and $\mu_{B}$ as $\mu_{S}\triangleq S(\bm{X}^{\prime})_{y_{t}}$ and $\mu_{B}\triangleq S(\bm{X})_{y_{t}}$, where $y_{t}$ is the target label. The $\Delta\mu$ is calculated as $\Delta\mu=\mu_{S}-\mu_{B}$. Given the null hypothesis $H_{0}:\mu_{B}+\tau=\mu_{S}$ ($H_{1}:\mu_{B}+\tau<\mu_{S}$) where $\tau\in[0,1]$, we claim that the suspicious model $S$ is stolen from the victim (with $\tau$-certainty) if and only if $H_{0}$ is rejected.
5.2 Main Results
As shown in Tables 4-5, Holmes is the only method that achieves reliable verification across all model stealing scenarios, outperforming baselines in most cases. For example, in the Soft Distill attack on the CIFAR-10 dataset, our method achieves a p-value of $10^{-8}$, significantly outperforming MOVE, the only baseline capable of successful verification in this scenario, which achieves a p-value of 0.01. In the cases of Direct-copy, Data Distill, and Fine-tuning, methods such as BadNets achieve competitive results. Nevertheless, our method can still easily make correct predictions in these cases. Besides, on the ImageNet dataset, our method demonstrates superior performance across all scenarios, including both model stealing cases and the independent model case. For example, in the Data Distill attack scenario, our method achieves a p-value of $10^{-20}$, outperforming DI’s $10^{-8}$, while all other baseline methods fail to provide model ownership verification. Additionally, in the independent model scenario for both CIFAR-10 and ImageNet datasets, our method consistently avoids misjudgments, while DI misclassifies the benign independent models as stolen models. This indicates that our method effectively decouples common features while emphasizing dataset-specific features, ensuring reliable model ownership verification across diverse datasets and model stealing attack methods.
5.3 Effects of Hyper-parameters